OpenStudy (anonymous):

In a local grade school, the ratio of boys to girls is 5: 7. There are 595 girls. How many boys are there?

OpenStudy (anonymous):

no of boys = 7*595/5=833

OpenStudy (katrinakaif):

Set up a proprotion: 5 x - = --- 7 595 Cross multiply: 7x = 595(5) x = 595(5)/7 and solve.

OpenStudy (anonymous):

Follow Katrina's steps. U will get answer as : 85 * 5 = 425